Balance the following reaction using oxidation numbers. б. х х, Нe Zn2+(aq) Vo2+ (aq) н,о() + H (aq) Zn(s) VO (aq)
Answer #1

1+ -> H+) + 2 vot & zn - voet + ant2 vot zn I noth ent? &- to x+2(-2) =+1/x=0 x+(-2)2+2). x-4=+1 x=+4 x=+5 oxidation half rea